The sum of the present ages of Eric nd his father is 58 years. In 10 years, his father will be twice as old as Eric will be at that time. Find their present ages.
Answer by checkley71(8403)

e+f=58 or e=58-f
(e+10)=f+10
2(58-f+10)=f+10
116-2f+20=f+10
-2f-f=10-20-116
-3f=126
f=-126/-3
f=42 age of the father 10 years ago.
e=58-42
e=16 age of Eric 10 years ago.
proof
42+16=58
also
2(58-42+10)=42+10
2*26=52
52=52